# Understanding the Training and Generalization of Pretrained Transformer for Sequential Decision Making

### v1

@@ E.1.1 Transformer for sequential decision making
E.1.2 Algorithm 1 implementation and pretraining details
@@ E.2 Environment Generation
Finite Pool of Environments. To study the behavior and test the performance of the transformer on finite possible environments (e.g., to see if and how $\texttt{TF}_{\hat{\theta}}$ converges to $\texttt{Alg}^{*}$), we also consider a finite pool of environments as the candidates of $\gamma$. Specifically, for some tasks we first sample finite environments (e.g., $\gamma_{1},\ldots,\gamma_{4}$) i.i.d. following the sampling rules previously described. We then set the environment distribution $\mathcal{P}_{\gamma}$ as a uniform distribution over the pool of sampled environments (e.g., $\{\gamma_{1},\ldots,\gamma_{4}\}$). We should notice this pool of environments does not restrict the context generation (if any): in both the pretraining and testing, contexts are generated following the rules previously described and are independent of the given finite pool.
@@ E.4.1 Details for Figure 1
In Figure 1 (a), we independently implement Algorithm 1 to train two transformer models on a dynamic pricing task, which has an infinite support of the prior environment distribution $\mathcal{P}_{\gamma}$ and $6$-dimensional contexts (more details can be found in Appendix E.2). For both models, the training parameters are identical except for the $M_{0}$ in Algorithm 1: the first curve (blue, thick line) uses $M_{0}=50$, while the second curve (orange, dashed line) uses $M_{0}=130$, meaning no transformer-generated data is utilized during training. All other parameters follow the configuration detailed in Appendix E.1. The figure shows the mean testing regret at each training iteration across 128 randomly sampled environments, with the shaded area representing the standard deviation. We further provide more results in Appendix C.1.

### v2

@@ E.1.1 Transformer for sequential decision making
E.1.2 Algorithm 1 implementation and pre-training details
@@ E.2 Environment Generation
Finite Pool of Environments. To study the behavior and test the performance of the transformer on finite possible environments (e.g., to see if and how $\texttt{TF}_{\hat{\theta}}$ converges to $\texttt{Alg}^{*}$), we also consider a finite pool of environments as the candidates of $\gamma$. Specifically, for some tasks we first sample finite environments (e.g., $\gamma_{1},\ldots,\gamma_{4}$) i.i.d. following the sampling rules previously described. We then set the environment distribution $\mathcal{P}_{\gamma}$ as a uniform distribution over the pool of sampled environments (e.g., $\{\gamma_{1},\ldots,\gamma_{4}\}$). We should notice this pool of environments does not restrict the context generation (if any): in both the pre-training and testing, contexts are generated following the rules previously described and are independent of the given finite pool.
@@ E.4.1 Details for Figure 2
In Figure 2 (a), we independently implement Algorithm 1 to train two transformer models on a dynamic pricing task, which has an infinite support of the prior environment distribution $\mathcal{P}_{\gamma}$ and $6$-dimensional contexts (more details can be found in Appendix E.2). For both models, the training parameters are identical except for the $M_{0}$ in Algorithm 1: the first curve (blue, thick line) uses $M_{0}=50$, while the second curve (orange, dashed line) uses $M_{0}=130$, meaning no transformer-generated data is utilized during training. All other parameters follow the configuration detailed in Appendix E.1. The figure shows the mean testing regret at each training iteration across 128 randomly sampled environments, with the shaded area representing the standard deviation. We further provide more results in Appendix C.1.2.